OpenStudy (anonymous):

Solve for x 120=1.5e^3x

OpenStudy (whpalmer4):

\[120=1.5e^{3x}\] Using the property of logs that \(\log a^n = n\log a\) take the log of both sides and solve for \(x\).

OpenStudy (whpalmer4):

(divide both sides by 1.5 first to isolate the exponential on the right)

OpenStudy (whpalmer4):

If you use the natural log, \(\ln e = 1\)...

OpenStudy (anonymous):

So is the answer 26.67 for x

OpenStudy (whpalmer4):

No, I don't think so...\(e^{80}\) is a big number!

OpenStudy (anonymous):

x=5.5406x10^34

OpenStudy (whpalmer4):

\[120 = 1.5e^{3x}\]Divide by 1.5 \[80 = e^{3x}\]Take natural log of both sides \[\ln 80 = \ln e^{3x} = 3x \ln e = 3x \] \[x=\frac{\ln 80}{3} \approx 4.932\]

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

so the x=4.932

OpenStudy (whpalmer4):

\[\ln 80 \approx 4.382\] \[\frac{\ln 80}{3} \approx 1.461\]

OpenStudy (anonymous):

ok I got that so then x=1.461

OpenStudy (whpalmer4):

Must have fat-fingered the calculation the first time, that's why I always check my answers!

OpenStudy (anonymous):

ok thanks lol

OpenStudy (whpalmer4):

yes, x = 1.461. \[120 = 1.5*e^{(3*1.461) }= 1.5*e^{4.383} = 1.5*80.07 \approx 120\] (answer is actually a smidge less than 1.461, clearly)